OpenStudy (anonymous):

is the system geometric? 2, -4,-16, -36

OpenStudy (anonymous):

identify the common ratio..

OpenStudy (anonymous):

2, -2, 3 , or no are the optons, i believe it is -2

OpenStudy (freckles):

is it geometric? well all all these quotients the same? -4/2=? -16/-4=? -36/-16=?

OpenStudy (anonymous):

i know they are all 2..

OpenStudy (anonymous):

would the signs stay the same though?

OpenStudy (freckles):

well the first one isn't 2 it is actually -2 for the first one

OpenStudy (freckles):

and 16/4 is 4 not 2

OpenStudy (freckles):

and 36/16 doesn't have a quotient of 2

OpenStudy (freckles):

all the quotients weren't the same so this is not a geometric seqeunce all the quotients have to be the same

OpenStudy (freckles):

I'm talking about the quotients of a term and its previous term of course

OpenStudy (anonymous):

wow i dont know what i was doing i was doing square for some reason.

OpenStudy (anonymous):

i thought you had to see which one would mutiply to fit them all but 16+16 is 32..

OpenStudy (anonymous):

so yeah itd be none.. thank you! i see

OpenStudy (freckles):

yea you can figure out the common ratio by saying what number times this term will give be the next term or you can figure that same common ratio number by taking a term and dividing it by its previous term if the sequence is of course geometric for example 2,4,8,16,32,... is geometric because \[\frac{4}{2}=2 \\ \frac{8}{4}=2 \\ \frac{16}{8}=2 \\ \frac{32}{16}=2 \] the common ratio is 2 or yeah you can look at it the way you did 2*2=4 4*2=8 8*2=16 16*2=32

OpenStudy (anonymous):

I see thanks for explaining that haha :)

OpenStudy (freckles):

np